A cistern can be filled by an inlet pipe in 6 hours & emptied by an outlet pipe in 10 hours. After 3 hours the outlet pipe is closed. How much time to fill up the remaining part of the tank?
Answer by richwmiller(17219)

3/6-3/10+x/6=1
x = 24/5 =4.8